OpenStudy (anonymous):

f(x) = 4x2 + 5 Find the value of f(x) when x=2.

OpenStudy (anonymous):

is that 4x^2?

OpenStudy (anonymous):

you have to substitute the value 2 in x of the function of 4x^2 +5 :)

OpenStudy (anonymous):

f(x=2) = 4(2)^2 + 5 :)

OpenStudy (anonymous):

Yes it is sorry, typo.

OpenStudy (anonymous):

so then 4*2^2+5= 69

OpenStudy (anonymous):

check it again i got a different answer :)

OpenStudy (anonymous):

4*2=8*8=64+5=69

OpenStudy (anonymous):

I still have the same answer.

OpenStudy (anonymous):

\[4x^2 +5 = 4(2)^2 +5 = 4(4) +5 = 16 + 5 = 21\]

OpenStudy (anonymous):

Oh okay thank you. So would 21 be my answer?

OpenStudy (anonymous):

yes, no problem :D